OpenStudy (anonymous):

Kendra is flying a kite 35 feet in the air. The angle of depression from the kite to where Kendra holds the string is . If Kendra holds the string 3 feet off the ground, what is the distance from her hands to the kite? Round your answer to the nearest tenth. a.37.7 feet b.41.3 feet c.60.4 feet d.66.0 feet

OpenStudy (anonymous):

please anyone help me

OpenStudy (anonymous):

@tarrelD99

OpenStudy (anonymous):

@TerrelD99 ... sorry messed up the name the first time

OpenStudy (anonymous):

If Kendra holds the string 3 feet off the ground, the distance from her hands to the kite is: 60.4 feet ***=35-3 = 32 ft sin(32) =32/x x = 32/sin(32) =60.3865572736188208= 60.4feet

OpenStudy (anonymous):

thank you soooooo much

OpenStudy (anonymous):

yvw.....\(\Huge\ddot\smile\)